COURSE SYLLABUS

Psychology 593LH How to Write in (Quantitative) Psychology Fall, 2010

Course Organizer: Lawrence Hubert 433 Psychology Building Office Hours: Wednesday 1–3

Course meets: Tuesday, 3–5, 708 Psychology Building

This workshop will be devoted to developing skills in writing for a psychology audience generally, and a quantitative audience more specifically. There are multiple emphases:

A) the formalities of writing a manuscript according to APA style (using the new Sixth Edition of the APA Manual);

B) writing coherent and correct (if ever so humble) English prose, with an emphasis on the Strunk and White bible of usage, The Elements of Style;

C) learning to navigate the mechanics of a L^{A}TEX and TEX system and editor;

D) mastering the specifics involved in writing technical and quantitative material (using the Not So Short Introduction to L^{A}TEX 2ε); E) familiarity with the APA class file and the UIUC Thesis class file; F) an ability to produce and edit one’s graphics with Adobe Illustrator (or similar graphics editor). The last four sessions will be partly devoted to editing/revising/critiquing writing samples provided by the workshop members.

Major References for the Course:

Publication Manual of the American Psychological Association, Sixth Edition, 2010 (referred to as APAM below). [Available from Amazon.com for $23.00; from our bookstore, it is about $40.00]

William Strunk and E. B. White, The Elements of Style (2000) (referred to as S & W below). [I have a free pdf of this item that you can have.]

The Not So Short Introduction to L^{A}TEX 2ε (referred to as NSSI below). (A free pdf of this item is available.)

Robert J. Sternberg (Editor), Guide to Publishing in Psychology Journals. (I have three of these available in Room 434.)

Writing an Empirical Article (Daryl Bem)

Writing a Literature Review (Nancy Eisenberg)

Title and Abstracts: They Only Sound Unimportant (Robert Sternberg)

Introducing Your Research Report: Writing the Introduction (Philip Kendall, Jennifer Silk, and Brian Chu)

Theories and Hypotheses (Abraham Tesser) Writing Effectively about Design (Harry Reis) Doing Data Analyses and Writing Up Their Results: Selected

Tricks and Artifices (Elena Grigorenko) Results That Get Results: Telling a Good Story (Peter Salovey) What Does it All Mean? The Discussion (Robert Calfee) Documenting Your Scholarship: Citations and References (Ran-

dolph Smith)

Writing for Your Referees (Robert Sternberg) Reading Reviews, Suffering Rejection, and Advocating for Your

Paper (Melissa Warren)

Rewriting the Psychology Paper (Richard Wagner) Article Writing 101: A Crib Sheet of 50 Tips for the Final Exam

(Robert Sternberg)

Special Symposium Issue on Writing Articles for Psychological Bulletin (1995, 118, Number 2):

Daryl J. Bem: Writing a Review Article for Psychological Bulletin

Ray Hyman: How to Critique a Published Article Robert Rosenthal: Writing Meta-Analytic Reviews Scott Maxwell and David Cole: Tips for Writing (and Reading)

Methodological Articles

Special Issue of Perspectives on Psychological Science (Volume 4, Number 1, 2009):

Denise Park: Publishing in the Psychological Sciences: Enhancing Journal Impact While Decreasing Author Fatigue

Shelley Taylor: Publishing in Scientific Journals: We’re Not Just Talking to Ourselves Anymore

Jerry Suls and René Martin: The Air We Breathe: A Critical Look at Practices and Alternatives in the Peer-Review Process

Martin Safer and Rong Tang: The Psychology of Referencing in Psychology Journal Articles

Seth Schwartz and Byron Zamboanga: The Peer-Review and Editorial System: Ways to Fix Something That Might be Broken

Nora Newcombe and Mark Bouton: Masked Reviews Are Not Fairer Reviews

David Trafimow and Stephen Rice: What If Social Scientists Had Reviewed Great Scientific Works of the Past

Raymond Nickerson: Are Social Scientists Harder on Their Colleagues Than Physical Scientists Were on Theirs in the Past: Commentary on Trafimow & Rice (2009)

M. Lynne Cooper: Problems, Pitfalls, and Promise in the Peer-Review Process: Commentary on Trafimow & Rice (2009)

\underline{Tentative topic sequences and readings}:

August 24: Overview of the course – your responsibilities.

August 31: APAM – Chapter 1: Writing for the Behavioral and Social Sciences APAM – Chapter 2: Manuscript Structure and Content

September 7: APAM – Chapter 3: Writing Clearly and Concisely APAM – Chapter 4: The Mechanics of Style

September 14: APAM – Chapter 5: Displaying Results APAM – Chapter 6: Crediting Sources

September 21: APAM – Chapter 7: Reference Examples APAM – Chapter 8: The Publication Process

September 28: S & W – Forward; Introduction; Elementary Rules of Usage; Ele-

mentary Principles of Composition

October 5: S & W – A Few Matters of Form; Words and Expressions Com-

monly Misused; An Approach to Style; Afterward; Glossary October 12: Producing Beautiful Documents with TEX and L^{A}TEX (Hubert’s

proselytizing talk (yet, once again))

October 19: NSSI – Chapters 1, 2, and 3 Things You Need to Know Typesetting Text Typesetting Mathematical Formulae

October 26: NSSI – Chapters 4, 5, and 6 Specialities Producing Mathematical Graphics Customizing L^{A}TEX

November 2: The APA class file (apa.cls) The UIUC thesis class file (uiucthesis.cls)

November 9, 16, 30, and December 7

The Lab: Editing/revising some of our own material Graphics editing with Illustrator

Reading and class presentations with discussion of selected chapters from Robert Sternberg, Guide to Publishing in Psychology Journals or the special issues of Psychological Bulletin and Perspectives on Psychological Science